Find the solution sets for the following:
3x^2-2x-8/x-1 >= 0

Answer by stanbon(75887)

Find the solution sets for the following:
3x^2-2x-8/x-1 >= 0
Factor:
[(x-2)(3x+4)]/[x-1] >=0
--------
Zeroes are x=2 and x=-4/3
The domain excludes x=1
--------------------------
Draw a number line and plot the values x=-4/3, x=1, x=2
Pick a test point from each interval of the number line bordered
by these values of x.
---------
If x=-10 you get (-12*-26)/-11>=0; this is false; no solution in (-inf,-4/3)
-------------
If x=0 you get (-2*4/-1)>=0; this is true; solutions in [-4/3,1)
--------------
If x=3/2 you get ((-1/2) * (17/2)/(1/2))>=0; false so no solution in (1,2)
---------------
If x=3 you get (8*13/2)>=0 ; true so solutions in [2,inf)
===========
Final Solution:
[-4/3,1) U [2,inf)
=====================
Cheers,
Stan H.